OpenStudy (anonymous):

2/3n + 6 = 1/4n -3 PLZ HELP !!!

OpenStudy (anonymous):

\[\frac{ 2 }{3 }n + 6 = \frac{ 1 }{ 4 }n - 3\]

OpenStudy (mertsj):

1. Subtract 1/4 n from both sides. 2. Subtract 6 from both sides 3. Multiply both sides by 12/5